Homework (Ch 02) Beth is a skilled toy maker who is able to produce both cars and puzzles. She has 8 hours a day to produce toys. The following table output resulting from various possible combinations of her time. shows the daily Hours Producing Produced Choice (Cars) (Puzzles) (Cars) (Puzzles) 0. 16 19 20 On the following graph, use the blue points (circle symbol) to plot Beths initial production possibilities frontier (PPF) Initial PPF New PPF 15 10 CARS

Indal PPlF New PPE Suppose Beth is currently using combination D, producing one car per day. Her opportunity cost of producing a second car per day is per day Now, suppose Bath is currerty using combination С, producing two cars per dr. Her OCCorturitv cost ol producing a third car per dar is per day From the previous analysis, you can determine that as Beth increases her production of cars, her opportunity cost of producing one more car Suppose EBeth buys a new tool that enaóhes hew to produce twice as many cars per how as betore, Dut ir doeset afact her abhoty to produce puzzia Use the green points (triangie symbol) to plot her new PPF on the previous Because she can now make more cars per hour, Beths opportunity cost of producino puzzles is
Answer #1

Opportunity cost is the value of next best alternative foregone.

Opportunity cost of producing additional car is the units of puzzles sacrificed.

1) To produce second car, producer has to decrease production of puzzles from 19 to 16. So, opportunity cost = 19 - 16 = 3 puzzles

2) To produce third car, producer has to decrease production of puzzles from 16 to 11. So, opportunity cost = 16 - 11 = 5 puzzles

3) Opportunity cost increases

4) lower than